# On the Convergence Theory of Pipeline Gradient-based Analog In-memory Training

### v2

@@ II-C Synchronous and asynchronous pipelines
In (10), the outer product $\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ is not a valid gradient with respect to any weight $W^{(m)}$, which is referred to as partially stale gradient and makes the analysis more challenging.
The main challenge coming from the AIMC training is the weight drift [20]. Consider the case where $W_{k}^{(m)}$ is already a non-zero critical point of $f(W)$ but not the one of $f(W;\xi)$, i.e. ${\mathbb{E}}_{\xi\sim{\mathcal{P}}}[\nabla_{W^{(m)}}f(W_{k}^{(m)};\xi)]=0$ but $\nabla_{W^{(m)}}f(W_{k}^{(m)};\xi)\neq 0$ for some $\xi$. After one iteration, the weight $W_{k}^{(m)}$ drifts because of the sample randomness
${\mathbb{E}}[W_{k+1}^{(m)}]-W_{k}^{(m)}=$ $\ -\frac{\alpha}{\tau}{\mathbb{E}}[|\nabla_{W^{(m)}}f(W_{k}^{(m)};\xi)|]\odot W_{k}^{(m)}\neq 0$ (11)
@@ III-A Assumptions
All activation function $g^{(m)}(\,\cdot\,)$, $m\in[M]$, is (i) Lipschitz continuous, i.e. $\|g^{(m)}(z)-g^{(m)}(z^{\prime})\|\leq L_{g,0}\|z-z^{\prime}\|$; (ii) smooth, i.e. $\|\nabla g^{(m)}(z)-\nabla g^{(m)}(z^{\prime})\|\leq L_{g,1}\|z-z^{\prime}\|$; (iii) centered, i.e. $g(0)=0$.
The conditions (i) and (ii) in Assumption 3 hold for many activation functions like Sigmoid, Tanh or smoothed ReLU functions, while (iii) can be achieved by adding an offset mapping before the activation function if there is a point $x$ such that $g(x)=0$.
@@
The loss $\ell(\cdot,\cdot)$ is smooth with respective the first input, i.e. $\|\nabla_{1}\ell(y_{1},y)-\nabla_{1}\ell(y_{2},y)\|\leq L_{\ell,1}\|y_{1}-y_{2}\|,\forall y_{1},y_{2},y$; lower bounded, i.e. $\ell(y_{1},y_{2})\geq f^{*},\forall y_{1},y_{2}$.
Aside from the assumptions, to avoid saturation during training, following the setting in [20], the saturation degree is assumed to be bounded.
@@
There exists a positive constant $W_{\max}<\tau$ such that all the weight $W^{(m)}_{k}$, $m\in[M]$ is $\ell_{\infty}$-norm bounded, i.e., $\|W^{(m)}_{k}\|_{\infty}\leq W_{\max,\infty}<\tau,\forall k\in[K]$. The ratio $W_{\max,\infty}/\tau$ is termed saturation degree.
From the physical perspective, Assumption 5 requires that the weight can be represented by a conductance state in the dynamic range [20]. It is possible since the trainable weights always remain bounded in a small neighborhood of the initialization points during training [33, 34, 35], especially for overparameterized networks in digital training. Given $W_{\max,\infty}$ is sufficiently large to include this neighborhood, Assumption 5 holds. While it is possible to prove this assumption as in [36], we adopt it to focus on revealing key insights into the effect of pipeline training.
@@ III-B Iteration complexity of pipeline training
where $B$ is the number of micro-batches, and the constant $S$ is given by $S:=\frac{W_{\max,\infty}/\tau^{2}}{1-W_{\max,\infty}/\tau^{2}}$.
The proof of Theorem 1 is deferred to Appendix D. Analog-SGD-SP converges in the rate ${\mathcal{O}}(\sqrt{{\sigma^{2}}/({BK})})$, which matches the rate of digital SGD [37]. Besides, the asymmetric bias introduces an additional term $\sigma^{2}S$, which is independent of $K$ and hence referred to as asymptotic error, which also occurs in Analog-SGD-WOP [20], and can be mitigated by a well-designed AIMC accelerator or error-correction algorithms [20, 19, 22]. Next, we provide the iteration complexity of Analog-SGD-AP.
@@
$E_{K}\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{K}}\frac{1}{1-(1+u)W_{\max,\infty}^{2}/\tau^{2}}+\sigma^{2}S^{\prime}+{\mathcal{O}}\left(\frac{1+1/u}{K}\right)$ (12)
@@
The proof of Theorem 2 is deferred to Appendix E. Similar to Analog-SGD-WOP or Analog-SGD-SP, we have asymptotic error $\sigma^{2}S^{\prime}$, which vanishes for a well-designed AIMC accelerator or can be mitigated by error-correction algorithms [20, 19]. Moreover, since ${\mathcal{O}}\left(\frac{1+1/u}{K}\right)$ is not the dominant term, one can set a small $u$ so that Analog-SGD-AP achieves the same convergence rate as the Analog-SGD-WOP [20].
@@
As a side product, we also provide an enhanced convergence analysis for SGD with asynchronous pipeline training on digital accelerators, detailed in Theorem 3 in the Appendix F. Our analysis, based on a novel Lyapunov function, shows that asynchronous pipeline effects appear only in higher-order terms, which does not slow down digital SGD, and the analysis is tighter than [32].
@@ III-C Proof sketch of Theorem 2
In this subsection, we highlight the key steps of the proof towards Theorem 2.
(Step I) Smoothness of the objective. Before analyzing the training dynamic, we show that the objective is smooth. Recall that $\nabla_{W^{(m)}}f(W;\xi)=\delta^{(m)}\otimes x^{(m)}$, we first show that under the assumptions of DNN structures, both $\delta^{(m)}$ and $x^{(m)}$ are smooth. Using slightly abusing notations, we denote the forward and backward signals computed by $W$ as $x^{(m)}$ and $\delta^{(m)}$, respectively; and denote the ones computed by $\tilde{W}$} as ${\tilde{x}}^{(m)}$ and $\tilde{\delta}^{(m)}$, respectively. The following lemma show that for each layer $m$, if we regard the signals as functions of weights, they are smooth functions.
@@
Based on this lemma, we prove the objective $f(W;\xi)$ is smooth.
@@
Lemma 2 enables us to adapt the tools from [37] to obtain the sufficient descent of the function value. However, the bias from the asymmetric update and stale signals from the asynchronous pipelining accumulates layer by layer within the DNN and eventually lead to an extra term in the upper bound. Next step we quantify the error of the stale signals.
(Step II) Error of stale signals. Similar to Step I, we begin from analyzing the error of $x^{(m)}$ and $\delta^{(m)}$. However, the stale signal involves the weights at different iterations and introduces additional errors. Intuitively, the error of forward signal ${\tilde{x}}^{(m)}_{k}$ comes from the stale weight $W^{(m)}_{k-(M-m)}$ as shown by (8); while the error of the backward signal arises from the error of the forward signal. It motivated us to bound the error by the delay. Since the error between latest $W^{(m^{\prime})}_{k}$ and the stale weight $W^{(m^{\prime})}_{k-(M-m)}$ can be decomposed as
@@
$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]-f(W_{k})\lesssim$ $\ -\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}$ (21)
$\ +\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}-\frac{1}{4\alpha}{\mathcal{O}}\left(\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]\right)+C_{\psi}\mathbb{E}_{\xi_{k}}[\psi_{k}]$
@@
(Step IV) Construction of the Lyapunov function. To compensate for this term in equation (21), we design the following Lyapunov function.
@@
and $C_{\psi}$ is a constant. To achieve the dynamics of $\Psi_{k}$, we derive that
Let the index of the inner summation of the first term in the RHS of (24) begin from $k-(M-m^{\prime})$ instead of $k+1-(M-m^{\prime})$, we have
@@
$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}+M\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k}\|^{2}-\sum_{m^{\prime}=1}^{M}m^{\prime}\|\Delta^{(m^{\prime})}_{k-(M-m^{\prime})}\|^{2}$
@@
Plugging (25) into (24) yields $\Psi_{k+1}\leq\Psi_{k}-\psi_{k}+M\|W_{k+1}-W_{k}\|^{2}$, which gives a negative $\psi_{k}$. We then show that the Lyapunov functions have sufficient descent if $C_{\psi}$ is properly chosen, i.e.,
$\mathbb{E}_{\xi_{k}}[{\mathbb{V}}_{k+1}]-{\mathbb{V}}_{k}\lesssim$ $\ -\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}.$ (27)
@@
Figure 4: Training ResNet models on CIFAR10 dataset via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn). (Left) Accuracy-vs-Epoch reflects sample complexity. (Middle) Accuracy-vs-Clock cycle. (Right) Speedup of the proposed pipeline training methods on AIMC accelerators as the number of stages increases compared with the single machine training case (stage=1). We ignore the communication latency since the asynchronous pipeline does not introduce obvious extra latency.

### v3

@@ II-C Synchronous and asynchronous pipelines
In (II-C), the outer product $\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ is not a valid gradient with respect to any weight $W^{(m)}$, which is referred to as partially stale gradient and makes the analysis more challenging.
The main challenge coming from the AIMC training is the weight drift [wu2024towards]. Consider the case where $W_{k}^{(m)}$ is already a non-zero critical point of $f(W)$ but not the one of $f(W;\xi)$, i.e. ${\mathbb{E}}_{\xi\sim{\mathcal{P}}}[\nabla_{W^{(m)}}f(W_{k}^{(m)};\xi)]=0$ but $\nabla_{W^{(m)}}f(W_{k}^{(m)};\xi)\neq 0$ for some $\xi$. After one iteration, the weight $W_{k}^{(m)}$ drifts because of the sample randomness
${\mathbb{E}}[W_{k+1}^{(m)}]-W_{k}^{(m)}=$ $\ -\frac{\alpha}{\tau}{\mathbb{E}}[|\nabla_{W^{(m)}}f(W_{k}^{(m)};\xi)|]\odot W_{k}^{(m)}$
$\neq$ $\ 0$ (11)
@@
Remark 2 (Communication overhead).
In all three pipeline strategies considered in this paper, inter-accelerator communication consists of passing activation vectors $x^{(m)}$ and backward signals $\delta^{(m)}$ between adjacent stages. The per-stage communication cost remains O(D) as M grows. Accordingly, deeper pipelines do not increase the asymptotic per-stage communication burden, although the realized wall-clock benefit still depends on bandwidth and latency remaining secondary to per-stage computation. Furthermore, the communication overhead can be reduced by overlapping gradient communication with computation as proposed in [chen2022sapipe, li2018pipe], thereby hiding inter-stage communication latency within the computation time of each clock cycle.
@@ III-A Assumptions
All activation functions are (i) Lipschitz continuous, i.e. $\|g^{(m)}(z)-g^{(m)}(z^{\prime})\|\leq L_{g,0}\|z-z^{\prime}\|$; (ii) smooth, i.e. $\|\nabla g^{(m)}(z)-\nabla g^{(m)}(z^{\prime})\|\leq L_{g,1}\|z-z^{\prime}\|$; (iii) centered, i.e. $g(0)=0$.
The conditions (i) and (ii) in Assumption 3 hold for many activation functions like Sigmoid, Tanh, or smoothed ReLU functions, while (iii) can be achieved by adding an offset mapping before the activation function if there is a point $x$ such that $g(x)=0$.
Remark 3 (Non-smooth activation functions).
Assumption 3 requires the activation functions to be smooth. For non-smooth activations such as ReLU, the signal smoothness (see Lemmas 1, 2 later) cannot be guaranteed in general. In this case, one can establish that the functions are semi-smooth [zou2019improved]. In the overparameterized regime, such a semi-smoothness property can be established by showing that the training trajectory remains within a small neighborhood of initialization where the set of active neurons is nearly fixed. Extending the convergence analysis of Analog-SGD-AP to non-smooth activations within this semi-smoothness framework is an interesting direction we leave for future work.
@@
The loss $\ell(\cdot,\cdot)$ is: (i) smooth with respect to the first input, i.e., $\|\nabla_{1}\ell(y_{1},y)-\nabla_{1}\ell(y_{2},y)\|\leq L_{\ell,1}\|y_{1}-y_{2}\|$; (ii) lower bounded, i.e., $\ell(y_{1},y_{2})\geq f^{*}$.
Aside from these assumptions, to avoid saturation during training, following the setting in [wu2024towards], the saturation degree is assumed to be bounded.
@@
There exists a positive constant $W_{\max,\infty}<\tau$ such that all the weight $W^{(m)}_{k}$, $m\in[M]$ is $\ell_{\infty}$-norm bounded, i.e., $\|W^{(m)}_{k}\|_{\infty}\leq W_{\max,\infty}<\tau,\forall k\in[K]$. The ratio $W_{\max,\infty}/\tau=\Theta(1)<1$ is termed saturation degree.
From the physical perspective, Assumption 5 requires that the weight can be represented by a conductance state in the dynamic range $(-\tau,\tau)$. Actually, the analog update dynamic induced by the hardware naturally constrains the weight, as shown by the following theorem.
Proposition 1 ([wu2024towards, Theorem 1]).
Given $\|W_{0}\|_{\infty}<\tau$ and for any sequence $\{\Delta W_{k}:k\in{\mathbb{N}}\}$, which satisfies $\|\Delta W_{k}\|_{\infty}\leq\tau$, it holds for any $W_{k}$ following (1) that $\|W_{k}\|_{\infty}<\tau,\,\forall k\in{\mathbb{N}}$.
Proposition 1 shows that if the magnitude of the update increment $\Delta W_{k}$ is bounded by $\tau$, then the entire trajectory remains within the dynamic range $(-\tau,\tau)$. We further establish a stronger result in Lemma 8: with a properly chosen stepsize $\alpha$, the Analog SGD update with asynchronous pipeline in (II-C) guarantees that the trajectory of $W_{k}$ stays within the dynamic range $(-\tau,\tau)$, without requiring any additional assumption on bounded update increments or bounded delay effects. However, we note that both Proposition 1 and Lemma 8 are established only for the finite-time setting, so that they do not rule out the extreme limit case where $\lim_{k\to\infty}|W_{k}|=\tau$. This is because the proof is built upon a worst-case analysis that accounts only for the gradient magnitude, rather than its direction. Under additional over-parameterization assumptions for DNNs [oymak2019overparameterized, liu2022loss, zou2019improved], it is possible to rule out this extreme case. However, a formal derivation is beyond the scope of this paper, whose primary focus is on the impact of pipeline parallelism. Moreover, in practice, the ratio $W_{\max,\infty}/\tau$ is much smaller than $1$. We empirically monitor the saturation degree, defined as $\|W_{k}^{(m)}\|_{\infty}/\tau$, throughout the training process. As illustrated in Section IV, the saturation degree for all layers remains remarkably stable and stays well below 10% across the full trajectory. Therefore, Assumption 5 is sensible in practice.
@@ III-B Iteration complexity of pipeline training
where $B$ is the number of micro-batches, and the constant $S$ is given by $S:=\frac{W_{\max,\infty}^{2}/\tau^{2}}{1-W_{\max,\infty}^{2}/\tau^{2}}$.
Analog-SGD-SP converges in the rate ${\mathcal{O}}(\sqrt{{\sigma^{2}}/({BK})})$, which matches the rate of digital SGD [bottou2018optimization]. Besides, the asymmetric bias introduces an additional term $\sigma^{2}S$, which is independent of $K$ and hence referred to as asymptotic error, which also occurs in Analog-SGD-WOP [wu2024towards], and can be mitigated by a well-designed AIMC accelerator or error-correction algorithms [wu2024towards, gokmen2020, gokmen2021]. Next, we provide the iteration complexity of Analog-SGD-AP.
@@
$E_{K}\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{K}}\frac{1}{1-(1+u)W_{\max,\infty}^{2}/\tau^{2}}$
$\ +\sigma^{2}S^{\prime}+{\mathcal{O}}\left(\frac{M^{3}(1+1/u)}{K}\right)$ (12)
@@
Similar to Analog-SGD-WOP or Analog-SGD-SP, we have asymptotic error $\sigma^{2}S^{\prime}$, which can be mitigated by error-correction algorithms [wu2024towards, gokmen2020]. Moreover, since ${\mathcal{O}}\left(\frac{1+1/u}{K}\right)$ is not the dominant term, one can set a small $u$ so that Analog-SGD-AP achieves the same convergence order as the Analog-SGD-WOP [wu2024towards] owing to our proof techniques in Section III-C.
@@
As a side product, we also provide an enhanced convergence analysis for SGD with asynchronous pipeline training on digital accelerators, detailed in Appendix F. Our analysis, based on a novel Lyapunov function, shows that asynchronous pipeline effects appear only in higher-order terms, which does not slow down digital SGD, and the analysis is tighter than [xu2020acceleration].
@@ III-C Proof sketch of Theorem 2
In this subsection, we highlight the key steps of the proof towards Theorem 2. We only summarize the key lemmas used for analysis. The proofs for each lemma can be found in E.
(Step I) Smoothness of the objective. Before analyzing the training dynamic, we show that the objective is smooth. Recall that $\nabla_{W^{(m)}}f(W;\xi)=\delta^{(m)}\otimes x^{(m)}$, we first show that under the assumptions of DNN structures, both $\delta^{(m)}$ and $x^{(m)}$ are smooth. Using slightly abusing notations, we denote the forward and backward signals computed by $W$ as $x^{(m)}$ and $\delta^{(m)}$, respectively; and denote the ones computed by $\tilde{W}$} as ${\tilde{x}}^{(m)}$ and $\tilde{\delta}^{(m)}$, respectively. The following lemma shows that for each layer $m$, if we regard the signals as functions of weights, they are smooth functions.
@@
Based on this lemma, we prove the objective is smooth.
@@
Lemma 2 enables us to adapt the tools from [bottou2018optimization] to obtain sufficient descent of the function value. However, the bias from the asymmetric update and stale signals from asynchronous pipelining accumulate layer by layer within the DNN and eventually lead to an extra term in the upper bound. Next step, we quantify the error of the stale signals.
(Step II) Error of stale signals. Similar to Step I, we begin from analyzing the error of $x^{(m)}$ and $\delta^{(m)}$. However, the stale signal involves the weights at different iterations and introduces additional errors. Intuitively, the error of the forward signal ${\tilde{x}}^{(m)}_{k}$ comes from the stale weight $W^{(m)}_{k-(M-m)}$ as shown by (8); while the error of the backward signal arises from the error of the forward signal. It motivated us to bound the error by the delay. Since the error between latest $W^{(m^{\prime})}_{k}$ and the stale weight $W^{(m^{\prime})}_{k-(M-m)}$ can be decomposed as
@@
$\ \mathbb{E}_{\xi_{k}}[f(W_{k+1})]-f(W_{k})$ (21)
$\lesssim$ $\ -\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}$
$\ +\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}-\frac{1}{4\alpha}{\mathcal{O}}\left(\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]\right)$
$\ +C_{\psi}\mathbb{E}_{\xi_{k}}[\psi_{k}]$
@@
(Step IV) Construction of the Lyapunov function. To address the key challenge that the accumulated delay error $\psi_{k}$ appears in the descent lemma (Lemma 4) as a positive term that prevents a clean telescoping argument, we augment the objective $f(W_{k})$ with an auxiliary function $\Psi_{k}$ that satisfies two properties: (i) $\Psi_{k+1}-\Psi_{k}$ produces a term that cancels $\psi_{k}$ in the descent of $f(W_{k})$ while other terms can be also canceled; and (ii) $\Psi_{k}$ itself is non-negative so that it works as an upper bound of $f(W_{k})$. Following this, construct a Lyapunov function as
@@
and $C_{\psi}$ is a constant. The weighting scheme $(k^{\prime}-k+M)$ in equation (23) is specifically designed to satisfy property (i). To see this, we view $\Psi_{k}$ as a weighted version of $\psi_{k}$, which yields a telescoping structure: it assigns larger weights to more recent iterates, reflecting that recent weight updates contribute more to the current staleness error. To be specific, the closer $k^{\prime}$ is to the current iteration $k$, the larger the effect of $\Delta_{k^{\prime}}^{(m^{\prime})}$ on the current iteration; if $k^{\prime}$ is farther from $k$, much of its delayed effect has already been consumed, so it receives a smaller weight. Enabled by this structure, we can prove the property of $\Psi_{k+1}\leq\Psi_{k}-\psi_{k}+M\left\|W_{k+1}-W_{k}\right\|^{2}$, which cancels the $\psi_{k}$ term in Lemma 4, while the extra $\left\|W_{k+1}-W_{k}\right\|^{2}$ term can be canceled by the $\frac{1}{4\alpha}{\mathcal{O}}\left(\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]\right)$ in (21).
To achieve the dynamics of $\Psi_{k}$, we derive that
$\Psi_{k+1}=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}(k^{\prime}-(k+1)+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (24)
$\ -\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
Let the index of the inner summation of the first term in the RHS of (III-C) begin from $k-(M-m^{\prime})$ instead of $k+1-(M-m^{\prime})$, we have
@@
$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
@@
$\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k}\|^{2}=$ $\ \sum_{m^{\prime}=1}^{M}\|W^{(m^{\prime})}_{k+1}-W^{(m^{\prime})}_{k}\|^{2}$ (26)
$=$ $\ \|W_{k+1}-W_{k}\|^{2}.$
Plugging (25) into (III-C) yields $\Psi_{k+1}\leq\Psi_{k}-\psi_{k}+M\|W_{k+1}-W_{k}\|^{2}$, which gives a negative $\psi_{k}$. We then show that the Lyapunov functions have sufficient descent if $C_{\psi}$ is properly chosen, i.e.,
$\ \mathbb{E}_{\xi_{k}}[{\mathbb{V}}_{k+1}]-{\mathbb{V}}_{k}$ (27)
$\lesssim$ $\ -\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}$
$\ +\alpha^{3}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{3}\sigma^{2}$
$\ +\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}.$ (28)
@@
Figure 3: ResNet10 @ CIFAR10 via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn). (Left) Accuracy-vs-Epoch reflects sample complexity. (Middle) Accuracy-vs-Clock cycle. (Right) Speedup of the proposed pipeline training methods on AIMC accelerators as the number of stages increases compared with the single machine training case (stage=1). We ignore the communication latency since the asynchronous pipeline does not introduce obvious extra latency.